### v1

@@ Abstract
Universal approximation property, neural networks, training algorithm, learning behavior
{MSCcodes}
41A30, 68T05, 68T07
@@ 1 Introduction
The universal approximation property (UAP) of neural networks is a cornerstone in the theoretical guarantee of deep learning, proving that even the simplest two-layer feedforward networks can approximate any continuous function [11, 21, 29]. This fascinating ability allows neural networks to replace critical, challenging components in existing frameworks to enhance efficiency [1, 23, 35, 47]. Despite the extensive study in various settings, existing UAP research rarely imposes restrictions on the network parameters [8, 13, 43]. However, in specific application cases, constraints are essential to meet certain requirements [26, 33].
As a constrained scenario, Qiu and Suda [34] empirically showed that, without altering the exact value of network weights, only permuting the initialized weights can achieve comparable or better performance for image classification tasks. This unique property makes the permutation-based training method attractive for specific hardware applications, and has been utilized as fixed-weight accelerators [26]. It can also facilitate the implementation of physical neural networks [14, 33]. Despite its impressive benefits in applications, the permutation training method suffers from an absence of theoretical guarantees regarding its effectiveness, which hinders further algorithmic and hardware development necessary to fully exploit its potential.
This paper establishes the first theoretical foundation of this method (to our best knowledge) by proving the UAP of a permutation-trained ReLU network with random initializations for any one-dimensional continuous function. Compared to the conventional UAP scenarios, the proof of permutation training encounters a significantly greater challenge, primarily due to the extreme constraints of maintaining the initialized weight values. The key proof idea is a four-pair construction of the step function approximators, which enables the approximation through a piecewise constant function [40]. Additionally, a reorganization method is proposed to eliminate the impact of the remaining weights.
Our numerical experiments not only validate our theoretical results by illustrating the widespread existence of the UAP of permutation training in diverse initializations, but also emphasize the effects of initializations on the permutation training performance. Moreover, the patterns observed during permutation training also highlight its potential in describing learning behavior, relating to topics like the pruning technique [16] and continual learning [31, 46]. Our main findings are summarized below:
@@ 1.1 Permutation training’s advantages in hardware implementation
Permutation training has been applied to training large models on extensive databases for image classification tasks [34], achieving performance that is comparable or even superior to conventional free-training training methods like Adam [25]. However, there is currently no evidence to report a significant advantage when applied to more diverse tasks on contemporary GPU-based hardware. Nevertheless, we believe it is highly suitable for the design of physical neural networks [33], since reconnecting the neurons is sometimes more convenient than altering the exact value. Therefore, permutation training may inspire alternative physical weight connection implementations, such as using fixed-weight devices controlled by a permutation circuit [34]. This idea has been applied to a fixed-weight network accelerator [26, 27].
Another potential application scenario is the physical neural networks with an explicit structure to store the weight value, such as the integrated photonic tensor core, a computing chip with specialized architecture [14]. This design has been successfully employed by international commercial companies in their photonic computing products. In each photonic tensor core, an array of phase-change cells are organized to separately store each element of the weight matrix, with their values adjusted through optical modulation of the transmission states of the cells. However, permutation training indicates that, in addition to changing the exact value, it is feasible to connect each cell with the permutation circuit for convenient reconnections. Therefore permutation training can facilitate the learning process.
@@ 1.2 Related works
The UAP has been extensively studied in various settings, leading to many efficient applications. It is well known that fully connected networks are universal approximators for continuous functions [11, 21, 29]. Further research primarily concentrated on how the width and depth of models influence their UAP. An estimation has been made of the lower bound for the minimum width required by networks to achieve UAP [4, 30]. It has also been demonstrated that the depth of the network is crucial, as increasing the network depth can enhance the expression power [39, 42, 44]. Our work considers the network with arbitrary width, aligning with conventional scenarios. However, the constraints on parameters imposed by permutation training introduce non-trivial complexities and challenges not encountered in traditional free training methods.
Permutation, as a typical group structure, has been systematically described [6]. In deep learning, it closely relates to permutation equivariant or invariant networks [10] designed to learn from symmetrical data [28, 45]. It is also evident in graph-structured data which inherently exhibit permutation invariance [32, 37]. However, permutation training is not limited to the issues with intrinsic symmetry.
As for the weight permutation attempts, Qiu and Suda [34] empirically proposed the first (to our knowledge) weight-permuted training method. This method preserves the initialized weight value, allowing more efficient and reconfigurable implementation of the physical neural networks [26, 27]. Our work provides theoretical guarantees of this method and considers some regression tasks numerically. Additionally, initialization can be improved by rewiring neurons from the perspective of computer networks [38], but the training methods are unchanged.
Permutation training is also closely related to the permutation symmetry and linear mode connectivity (LMC) [12, 17]. The LMC suggests that after a proper permutation, most SGD solutions under different initialization will fall in the same basin in the loss landscape. Similarly, permutation training also seeks a permutation to improve performance. Therefore, the search algorithms utilized in LMC indicate the possibility of more efficient permutation training algorithms, as the fastest algorithm can search a proper permutation of large ResNet models in seconds to minutes [2, 24].
@@ 2.1 Nerual networks architecture
We start with a one-hidden-layer feed-forward ReLU network with $N$ hidden neurons in even numbers (i.e., $N=2n$). It has the form of a linear combination of ReLU basis functions like
@@
where all parameters are scalars when approximating one-dimensional functions, i.e., $w_{i},b_{i},a_{i},c\in\mathbb{R}$. Since ReLU activation is positively homogeneous i.e., $\text{ReLU}(\lambda x)=\lambda\text{ReLU}(x)$ for all $\lambda>0$, we consider a homogeneous case with $w_{i}=\pm 1$, and divide the basis functions into two parts as
(1) $\phi_{i}^{\pm}(x)=\text{ReLU}\big{(}\pm(x-b_{i})\big{)},\quad i=1,2,...,n,$
where the biases $\{b_{i}\}_{i=1}^{n}$ determine the basis locations  [^1] In this paper, we use this notation $\{a_{i}\}_{i=1}^{n}:=\{a_{1},\cdots,a_{n}\}$ to represent a set with $n$ elements, while $(a_{i})_{i=1}^{n}:=(a_{1},\cdots,a_{n})$ for a vector with $n$ components.. Next, we introduce two factors $\alpha,\gamma$ to adjust the network’s output, leading to an additional one-dimensional linear layer. While this layer is not essential for achieving UAP, it does simplify the proof and offer practical value. The network’s output function $f^{\text{NN}}$ gives
(2) $f^{\text{NN}}(x)=\alpha+\gamma\sum_{i=1}^{n}\big{[}p_{i}\phi_{i}^{+}(x)+q_{i}\phi_{i}^{-}(x)\big{]},$
where $\theta^{(2n)}=(p_{1},q_{1},...,p_{n},q_{n})\in\mathbb{R}^{2n}$ are the coefficient vector of basis functions, which also correspond to the parameters in the second hidden layer of the network.
@@ 2.2 Permutation and corresponding properties
For a vector $V^{(n)}=(v_{1},v_{2},\cdots,v_{n})$, the permutation $\tau$ is a bijection from the index set $\{1,2,\cdots,n\}$ to itself, such that $\tau(i)=j$ implies $v_{\tau(i)}=v_{j}$.
We denote the permuted vector as $\tau(V^{(n)})$. Notice that all permutations of $V^{(n)}$ can form a group $S_{n}$, which is closed under composition. Precisely, for any $\pi,\tau\in S_{n}$, there is a $\rho\in S_{n}$ such that $\rho(i)=\pi(\tau(i))$ for all $i\in\{1,\cdots,n\}$. This property leads to the major advance of permutation training: ideally, the final weight $\theta_{T}$ can be achieved by permuting the initialized weight $\theta_{0}$ only once, i.e., there exists a $\tau\in S_{2n}$ such that $\theta_{T}=\tau(\theta_{0})$. This one-step nature significantly distinguishes permutation training from other iterative training methods like Adam.
@@ 2.3 Weight configuration and main theorems
Let $\alpha=0,\gamma=1$. For any function $f^{*}\in C([0,1])$ and any small number $\varepsilon>0$, there exists a large integer $n\in\mathbb{Z}^{+}$, for $f_{\text{pc}}^{\text{NN}}$ in Eq. (2) with equidistantly distributed $B^{(n)}=(b_{i})_{i=1}^{n}:=\left(0,\tfrac{1}{n-1},\cdots,1\right)$ and $W^{(2n)}=(\pm b_{i})_{i=1}^{n}:=(+b_{1},-b_{1},...,+b_{n},-b_{n})$, along with a permuted coefficients $\theta^{(2n)}=\tau(W^{(2n)})$ such that $|f_{\text{pc}}^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$.
@@
Given a probability threshold $\delta\in(0,1)$, for any function $f^{*}\in C([0,1])$ and any small number $\varepsilon>0$, there exists a large integer $n\in\mathbb{Z}^{+}$, and $\alpha,\gamma\in\mathbb{R}$ for $f_{\text{rand}}^{\text{NN}}$ in Eq. (2) with randomly initialized $B_{rand}^{(n)}\sim\mathcal{U}[0,1]^{n}$ and pairwisely randomly initialized $W_{rand}^{(2n)}=(\!\pm p_{i})_{i=1}^{n}$, $p_{i}\sim\mathcal{U}[0,1]$, along with a permuted coefficients $\theta^{(2n)}=\tau(W_{rand}^{(2n)})$, such that with probability $1-\delta$, $|f_{\text{rand}}^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$.
@@ 2.4 Proof ideas
To identify the UAP of our network Eq. (2) in $C([0,1])$, we employ a piecewise constant function, which is a widely-used continuous function approximator [40], and can be expressed as a summation of several step functions. Next, we demonstrate that our networks can approximate each step function. In this spirit, our constructive proof includes the following three steps (illustrated in Fig. 1):
@@
Step 1 can be achieved by dividing the range of $f^{*}$ with a uniform height of each step in $g$ (see Fig. 1(a)). The statement is outlined below.
@@
For any function $f^{*}\in C([0,1])$ and any small number $\varepsilon>0$, there is a piecewise constant function $g$ with a common jump $\Delta h\leq\varepsilon$, such that $|g(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$. Moreover, the function $g$ can be written as a summation of $J$ step functions $\{f_{s_{j}}\}_{j=1}^{J}$ as the following form,
(3) $g(x)=\sum_{j=1}^{J}a_{j}f_{s_{j}}(x)=\sum_{j=1}^{J}a_{j}\Delta h\chi(x-s_{j}),\quad a_{j}=\pm 1,\>s_{j}\in[0,1],\>J\in\mathbb{Z}^{+}.$
@@
Proof 2.6.
The function $g$ can be constructed explicitly. According to the Stone-Weierstrass theorem [41], we assume $f^{*}$ to be a polynomial function for simplicity. Let the range of $f^{*}$ be covered by an interval $[c_{1}\Delta h,c_{2}\Delta h]$ with two integers $c_{1},c_{2}\in\mathbb{Z}$. Then for any $j=1,\cdots,J$, we choose $s_{j}\in[0,1]$ such that $f^{*}(s_{j})=(k_{j}+0.5)\Delta h$ for some $k_{j}\in\mathbb{Z}$. The step locations $s_{1}<\cdots<s_{J}$ are distinct from each other otherwise there will be a discontinuity. The step sign $a_{j}$ is determined according to the values of $f^{*}$ on $[s_{j-1},s_{j+1}]$. It is easy to verify such construction satisfies our requirements.
The execution of step 2 is inspired by the divide-and-conquer algorithm in computer science [20]. For each step function $f_{s_{j}}$ in $g$, we select basis functions to construct a step function approximator $f_{s_{j}}^{\text{NN}}$, then sum them up to approach $g$. This step-matching construction utilizes four pairs of basis functions {±$\{\pm b_{i}\}_{i=1}^{4}$ (shown in Fig. 1(b)), and establishing a one-to-one mapping between coefficients and biases, i.e., $\{p_{i},q_{i}\}_{i=1}^{4}=\{\pm b_{i}\}_{i=1}^{4}$. It ensures that each coefficient and location are uniquely assigned and prevents conflict between different approximators.

### v2

@@ Abstract
Universal approximation property , neural networks , training algorithm , learning behavior.
MSC:
[2020] 41A30 , 68T05 , 68T07
journal: Neural Networks
[…]
[BNU]organization=School of Mathematical Sciences, Laboratory of Mathematics and Complex Systems, MOE, Beijing Normal University,city=Beijing, postcode=100875, country=China
[…]
[PolyU1]organization=Department of Applied Mathematics, The Hong Kong Polytechnic University,addressline=Hung Hom, Kowloon, country=Hong Kong
[…]
[PolyU2]organization=Department of Applied Mathematics & Research Institute for Smart Energy, The Hong Kong Polytechnic University,addressline=Hung Hom, Kowloon, country=Hong Kong
@@ 1 Introduction
The Universal Approximation Property (UAP) of neural networks is a cornerstone in the theoretical guarantee of deep learning, proving that even the simplest two-layer feedforward networks can approximate any continuous function (Cybenko, 1989; Hornik et al., 1989; Leshno et al., 1993). This fascinating ability allows neural networks to replace critical, challenging components in existing frameworks to enhance efficiency (Raissi et al., 2019; Lu et al., 2021). Despite the extensive study in various settings (Castro et al., 2000; Fan et al., 2020; Wang and Qu, 2022), existing UAP research rarely imposes restrictions on the network parameters. However, in specific application cases, constraints are essential to meet certain requirements (Nugent, 2005; Kosuge et al., 2021a).
As a constrained scenario, Qiu and Suda (2020) empirically showed that, without altering the exact value of network weights, only permuting the initialized weights can achieve comparable or better performance for image classification tasks. This unique property makes the permutation-based training method attractive for specific hardware applications, and has been utilized as fixed-weight accelerators (Kosuge et al., 2021a). It can also facilitate the implementation of physical neural networks (Nugent, 2005; Feldmann et al., 2021). Despite its impressive benefits in applications, the permutation training method suffers from an absence of theoretical guarantees regarding its effectiveness, which hinders further algorithmic and hardware development necessary to fully exploit its potential.
This paper establishes the first theoretical foundation of this method (to our best knowledge) by proving the UAP of a permutation-trained Rectified Linear Unit (ReLU) network with random initializations for any one-dimensional continuous function. Compared to the conventional UAP scenarios, the proof of permutation training encounters a significantly greater challenge, primarily due to the extreme constraints of maintaining the initialized weight values. The key proof idea is a four-pair construction of the step function approximators, which enables the approximation through a piecewise constant function (Stein and Shakarchi, 2009). Additionally, a reorganization method is proposed to eliminate the impact of the remaining weights.
Our numerical experiments not only validate our theoretical results by illustrating the widespread existence of the UAP of permutation training in diverse initializations, but also emphasize the effects of initializations on the permutation training performance. Moreover, the patterns observed during permutation training also highlight its potential in describing learning behavior, relating to topics like the pruning technique (Frankle and Carbin, 2019) and continual learning (Maltoni and Lomonaco, 2019; Zeng et al., 2019). Our main findings are summarized below:
@@ 1.1 Permutation training’s advantages in hardware implementation
Permutation training has been applied to training large models on extensive databases for image classification tasks (Qiu and Suda, 2020), achieving performance that is comparable or even superior to conventional free-training training methods like Adam (Kingma and Ba, 2015). However, there is currently no evidence to report a significant advantage when applied to more diverse tasks on contemporary Graphics Processing Unit (GPU)-based hardware. Nevertheless, we believe it is highly suitable for the design of physical neural networks (Nugent, 2005), since reconnecting the neurons is sometimes more convenient than altering the exact value. Therefore, permutation training may inspire alternative physical weight connection implementations, such as using fixed-weight devices controlled by a permutation circuit (Qiu and Suda, 2020). This idea has been applied to a fixed-weight network accelerator (Kosuge et al., 2021a, b).
Another potential application scenario is the physical neural networks with an explicit structure to store the weight value, such as the integrated photonic tensor core, a computing chip with specialized architecture (Feldmann et al., 2021). This design has been successfully employed by international commercial companies in their photonic computing products. In each photonic tensor core, an array of phase-change cells are organized to separately store each element of the weight matrix, with their values adjusted through optical modulation of the transmission states of the cells. However, permutation training indicates that, in addition to changing the exact value, it is feasible to connect each cell with the permutation circuit for convenient reconnections. Therefore permutation training can facilitate the learning process.
@@ 1.2 Related works
The UAP has been extensively studied in various settings, leading to many efficient applications. It is well known that fully connected networks are universal approximators for continuous functions (Cybenko, 1989; Hornik et al., 1989; Leshno et al., 1993). Further research primarily concentrated on how the width and depth of models influence their UAP. An estimation has been made of the lower bound for the minimum width required by networks to achieve UAP (Cai, 2023; Lu et al., 2017). It has also been demonstrated that the depth of the network is crucial, as increasing the network depth can enhance the expression power (Shen et al., 2022; Telgarsky, 2016; Yarotsky, 2017). Our work considers the network with arbitrary width, aligning with conventional scenarios. However, the constraints on parameters imposed by permutation training introduce non-trivial complexities and challenges not encountered in traditional free training methods.
Permutation, as a typical group structure, has been systematically described (Cameron, 1999). In deep learning, it closely relates to permutation equivariant or invariant networks (Cohen and Welling, 2016) designed to learn from symmetrical data (Lee et al., 2019; Zaheer et al., 2017). It is also evident in graph-structured data which inherently exhibit permutation invariance (Maron et al., 2019; Satorras et al., 2021). However, permutation training is not limited to the issues with intrinsic symmetry.
As for the weight permutation attempts, Qiu and Suda (2020) empirically proposed the first (to our knowledge) weight-permuted training method. This method preserves the initialized weight value, allowing more efficient and reconfigurable implementation of the physical neural networks (Kosuge et al., 2021a, b). Our work provides theoretical guarantees of this method and considers some regression tasks numerically. Additionally, initialization can be improved by rewiring neurons from the perspective of computer networks (Scabini et al., 2022), but the training methods are unchanged.
Permutation training is also closely related to the permutation symmetry and Linear Mode Connectivity (LMC) (Entezari et al., 2021; Frankle et al., 2020). The LMC suggests that after a proper permutation, most stochastic gradient descent solutions under different initialization will fall in the same basin in the loss landscape. Similarly, permutation training also seeks a permutation to improve performance. Therefore, the search algorithms utilized in LMC indicate the possibility of more efficient permutation training algorithms, as the fastest algorithm can search a proper permutation of large ResNet models in seconds to minutes (Ainsworth et al., 2023; Jordan et al., 2023).
@@ 2.1 Nerual networks architecture
We start with a one-hidden-layer feed-forward ReLU network with $N$ hidden neurons. It has the form of a linear combination of ReLU basis functions like
@@
where all parameters are scalars when approximating one-dimensional functions, i.e., $w_{i},b_{i},a_{i},c\in\mathbb{R}$. Since ReLU activation is positively homogeneous i.e., $\text{ReLU}(\lambda x)=\lambda\text{ReLU}(x)$ for all $\lambda>0$, we consider a homogeneous case with $w_{i}=\pm 1$. To facilitate our construction below, we assume $N$ is an even number (i.e., $N=2n$) and the basis functions located pairwisely as
$\phi_{k}^{\pm}(x)=\text{ReLU}\big{(}\pm(x-b_{k})\big{)},\quad k=1,2,...,n,$ (1)
where the biases $\{b_{k}\}_{k=1}^{n}$ determine the basis locations  [^2] In this paper, we use this notation $\{a_{k}\}_{k=1}^{n}:=\{a_{1},\cdots,a_{n}\}$ to represent a set with $n$ elements, while $(a_{k})_{k=1}^{n}:=(a_{1},\cdots,a_{n})$ for a vector with $n$ components.. The requirement of even $N$ can be removed by adding a ”ghost basis function” $\phi_{k}^{\pm}(x)$ with $a_{i}=0$ if it can be paired with an existing $\phi_{k}^{\mp}(x)$ also with $a_{i}=0$. Nevertheless, we retain this condition for simplicity. Next, we introduce two factors $\alpha,\gamma$ to adjust the network’s output, leading to an additional one-dimensional linear layer. While this layer is not essential for achieving UAP, it does simplify the proof and offer practical value. The network’s output function $f^{\text{NN}}$ gives
$f^{\text{NN}}(x)=\alpha+\gamma\sum_{k=1}^{n}\big{[}p_{k}\phi_{k}^{+}(x)+q_{k}\phi_{k}^{-}(x)\big{]},$ (2)
where $\theta^{(2n)}=(p_{1},q_{1},...,p_{n},q_{n})\in[-1,1]^{2n}$ are the coefficient vector of basis functions, which also correspond to the parameters in the second hidden layer of the network.
@@ 2.2 Permutation and corresponding properties
For a vector $V^{(m)}=(v_{1},v_{2},\cdots,v_{m})$, the permutation $\tau$ is a bijection from the element set $\{v_{i}\}_{i=1}^{m}$ to itself.
We denote the permuted vector as $\tau(V^{(m)})=(\tau(v_{i}))_{i=1}^{m}$. Notice that all permutations of $V^{(m)}$ can form a group $S_{m}$, which is closed under composition. Precisely, for any $\pi,\tau\in S_{m}$, there is a $\rho\in S_{m}$ such that $\rho(v_{i})=\pi(\tau(v_{i}))$ for all $v_{i}$ within $V^{(m)}$. This property leads to the major advance of permutation training: ideally, the final weight $\theta_{T}$ can be achieved by permuting the initialized weight $\theta_{0}$ only once, i.e., there exists a $\tau\in S_{2n}$ such that $\theta_{T}=\tau(\theta_{0})$. This one-step nature significantly distinguishes permutation training from other iterative training methods like Adam.
@@ 2.3 Weight configuration and main theorems
Let $\alpha=0,\gamma=1$. For any function $f^{*}\in C([0,1])$ and any small number $\varepsilon>0$, there exists a large integer $n\in\mathbb{Z}^{+}$, for $f^{\text{NN}}$ in Eq. (2) with equidistantly distributed $B^{(n)}=(b_{i})_{i=1}^{n}:=\left(0,\tfrac{1}{n-1},\cdots,1\right)$ and $W^{(2n)}=(\pm b_{i})_{i=1}^{n}:=(+b_{1},-b_{1},...,+b_{n},-b_{n})$, along with a permuted coefficients $\theta^{(2n)}=\tau(W^{(2n)})$ such that $|f^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$.
@@
Given a probability threshold $\delta\in(0,1)$, for any function $f^{*}\in C([0,1])$ and any small number $\varepsilon>0$, there exists a large integer $n\in\mathbb{Z}^{+}$, and $\alpha,\gamma\in\mathbb{R}$ for $f^{\text{NN}}$ in Eq. (2) with randomly initialized $B_{rand}^{(n)}\sim\mathcal{U}[0,1]^{n}$ and pairwisely randomly initialized $W_{rand}^{(2n)}=(\!\pm p_{i})_{i=1}^{n}$, $p_{i}\sim\mathcal{U}[0,1]$, along with a permuted coefficients $\theta^{(2n)}=\tau(W_{rand}^{(2n)})$, such that with probability $1-\delta$, $|f^{\text{NN}}(x)-f^{*}(x)|\leq\varepsilon$ for all $x\in[0,1]$.
@@ 2.4 Proof ideas
To identify the UAP of our network Eq. (2) in $C([0,1])$, we employ a piecewise constant function, which is a widely-used continuous function approximator (Stein and Shakarchi, 2009), and can be expressed as a summation of several step functions. Next, we demonstrate that our networks can approximate each step function. In this spirit, our constructive proof includes the following three steps (illustrated in Fig. 1):
@@
Step 1 can be achieved by cutting the range of $f^{*}$ into subregions with an equal width $\Delta h$, and then constructing a step function at each point where $f^{*}$ crosses a boundary of these subregions, leading to an approximator with an error $\Delta h$ (illustrated in Fig. 1(a)). The statement is outlined below.
@@
For any function $f^{*}\in C([0,1])$ and any small number $\varepsilon^{\prime}>0$, there is a piecewise constant function $g$ with a common jump $\Delta h\leq\varepsilon^{\prime}$, such that $|g(x)-f^{*}(x)|\leq\varepsilon^{\prime}$ for all $x\in[0,1]$. Moreover, the function $g$ can be written as a summation of $J$ step functions $\{f_{s_{j}}\}_{j=1}^{J}$ as the following form,
$g(x)=\sum_{j=1}^{J}a_{j}f_{s_{j}}(x)=\sum_{j=1}^{J}a_{j}\Delta h\chi(x-s_{j}),\quad a_{j}=\pm 1,\>s_{j}\in[0,1],\>J\in\mathbb{Z}^{+},$ (3)
@@
Proof.
The function $g$ can be constructed explicitly. According to the Stone-Weierstrass theorem (Stone, 1948), we assume $f^{*}$ to be a polynomial function for simplicity. Let the range of $f^{*}$ be covered by an interval $[k_{\text{min}}\Delta h,k_{\text{max}}\Delta h]$ with two integers $k_{\text{min}},k_{\text{max}}\in\mathbb{Z}$. Then denote $J$ as the number of intersections between $f^{*}$ and parallel lines $y=(k+0.5)\Delta h$, $k=k_{\text{min}},k_{\text{min}}+1,\cdots,k_{\text{max}}-1$. Since $f^{*}$ is a polynomial function, $J$ can be assumed to be finite (Otherwise $f^{*}$ will be a constant function due to the fundamental theorem of algebra (Hungerford, 2012), making it easily approximated by $g$). The scenario with $J=0$ is also trivial since it indicates that $f^{*}$ lies in $[(k_{0}-0.5)\Delta h,(k_{0}+0.5)\Delta h]$ for some integer $k_{0}\in[k_{\text{min}},k_{\text{max}}]$, such that $f^{*}$ can be approached by the constant function $y=k_{0}\Delta h$ with an error $\Delta h$.
Hence, for any $j=1,\cdots,J$, we choose $s_{j}\in[0,1]$ such that $f^{*}(s_{j})=(k_{j}+0.5)\Delta h$ for some $k_{j}\in\mathbb{Z}$. The step locations $s_{1}<\cdots<s_{J}$ are distinct since any repetition would contradict the continuity of the target function $f^{*}$. The step sign $a_{j}$ is determined according to the values of $f^{*}$ on $[s_{j-1},s_{j+1}]$. It is easy to verify such construction satisfies our requirements. ∎
The execution of step 2 is inspired by the divide-and-conquer algorithm in computer science (Hopcroft et al., 1983). For each step function $f_{s_{j}}$ in $g$, we select basis functions to construct a step function approximator $f_{s_{j}}^{\text{NN}}$, then sum them up to approach $g$. This step-matching construction utilizes four pairs of basis functions {±$\{\pm b_{i}\}_{i=1}^{4}$ (shown in Fig. 1(b)), and establishing a one-to-one mapping between coefficients and biases, i.e., $\{p_{i},q_{i}\}_{i=1}^{4}=\{\pm b_{i}\}_{i=1}^{4}$. It ensures that each coefficient and location are uniquely assigned and prevents conflict between different approximators.